OpenStudy (anonymous):

Find the maximum and minimum values of the function on the given interval. y = x2 – 6x – 5, [-1, 0] I'm pretty sure the minimum is -5

zepdrix (zepdrix):

\[\huge y=x^2-6x-5\]\[\huge y'=2x-6\]We'll set this equal to zero to find critical points and also factor out the 2.\[\huge 0=2(x-3)\]

zepdrix (zepdrix):

Hmm, where'd you get the -5 from? :)

OpenStudy (anonymous):

yes... you'r correct...

zepdrix (zepdrix):

Oh i see, you mean after comparing all of the points ^^ yah good job mcd.